# Multi-Partite Output Regulation of Multi-Agent Systems

### v1

@@ III-A Distributed Control Law
Any linear time-invariant distributed controller solving the CORP for the MAS composed of (III-A) and (2) under an augmented unsigned directed graph solves the MORP under the same graph.
In the light of Lemma 2, we consider the following feedforward-based distributed control law proposed in [5]
@@
where $\eta_{i}\in\mathbb{R}^{n_{0}}$ is the estimate of the exogenous signal $v$. The state equation in (III-A) is called a distributed observer. Moreover, $\mu\in\mathbb{R}$, $K_{1i}\in\mathbb{R}^{m_{i}\times n_{i}}$, and $K_{2i}\in\mathbb{R}^{m_{i}\times n_{0}}$ are control parameters to be designed. Compared to the distributed observer in [5], the one in (III-A) uses the absolute value of the adjacency matrix entries to be readily applicable even if the given graph is signed.
@@ III-B MORP: Necessary and Sufficient Conditions
The augmented signed directed graph $\mathcal{G}(\bar{\mathcal{A}})$ contains a directed spanning tree.
@@
The $k$-partition transformation term $s_{i}$ is incorporated into (3) due to Lemma 2. Except for this term, Condition 3 is standard for the studies investigating the CORP with the feedforward approach (e.g., see [5, 9]). The articles [19, 17] studying the BORP with the feedforward approach consider the disturbance-free (i.e., $E_{i}=0$) and direct feedthrough-free (i.e., $D_{i}=0$) follower dynamics. However, introducing a gauge transformation term in the regulator equations for each follower can extend their distributed controllers to take the effect of disturbance and direct feedthrough into account. To avoid such modification, the authors in [18] assume that agents in different sets in partition are subject to disturbances that are additive inverses of each other. This can be impractical in real-world applications, for instance, consider networked UAVs operating in the same environment. Thus, this article does not make that assumption about disturbances.
@@ III-C Discussion: Synthesis of Control Parameters
This subsection presents the first design strategy for the control parameters and highlights an associated drawback. To this end, let $\mathcal{L}_{u}$ denote the Laplacian matrix of the unsigned directed graph $\mathcal{G(|\mathcal{A}|)}$ and $\mathcal{H}=\mathcal{L}_{u}+\mathcal{F}$. Let $\lambda_{i}(\mathcal{H})$ and $\lambda_{j}(A_{0})$ denote the eigenvalues of $\mathcal{H}$ and $A_{0}$ for $i=1,\ldots,N$, $j=1,\ldots,n_{0}$.
@@
    Select $\mu$ according to the inequality111 The sufficient and necessary conditions for the Hurwitzness of the matrix $A_{\mu}=(I_{N}\otimes A_{0})-\mu(\mathcal{H}\otimes I_{n_{0}})$, which is the system matrix of the distributed observer in compact form, are used in the proof of Theorem 1. These conditions are given in Lemma 4 of [31]. However, the lower bound on $\mu$ in (8), due to Lemma 3 given in Appendix, has no conservatism compared to the bound in Lemma 4 of [31]. Another lower bound is also provided in Lemma 3.2 of [1]. But, one can easily construct a counterexample to the first statement of that lemma by considering a Hurwitz $A_{0}$.
@@
Since $\mathrm{spec}(\mathcal{H})\subsetneq\mathrm{ORHP}$ under Condition 4 (e.g., see Lemma 1 in [5]), any positive $\mu$ satisfies the inequality (8) if $\mathrm{spec}(A_{0})\subsetneq\mathrm{CLHP}$. Therefore, the design of $\mu$ is independent of the eigenvalues of $\mathcal{H}$ (i.e., the spectral property of the graph $\mathcal{G}(\bar{\mathcal{A}})$) when the exosystem generates a linear combination of constant signals, sinusoidal signals, polynomial signals, for instance, ramp signals, and exponentially converging signals.
@@
Though the design based on Remark 6 is straightforward to apply, it has a drawback in that the pair $(X_{i},U_{i})$ for each follower depends on $s_{i}$. Therefore, whenever the $k$-partition transformation is updated, it necessitates each follower to recompute a solution pair for the regulator equations (3) unless $s_{i}$ remains unchanged. Hence, given a finite set of $k$-partition transformation terms with $M$ elements, each follower computes $M$ solution pairs to (3). This drawback can render the design strategy impractical in some applications. For instance, consider low-cost networked UAVs trying to avoid radars. In this scenario, the set of $k$-partition transformation terms for each follower may not be known before the operation. Therefore, it may be desirable in terms of computational cost to find a solution pair for the regulator equations (3) once and use it throughout the operation.
@@ IV MORP: Partition-Independent Solvability
This section first provides the partition-independent sufficient conditions for the solvability of the MORP. For each $i\in\mathcal{N}$, these conditions include the partition-independent regulator equations, as in the CORP with the feedforward approach, and an LME depending solely on $B_{i}$, $E_{i}$, and $D_{i}$. The constructive nature of the result yields the second design strategy that does not suffer from the drawback discussed in Remark 7. Then, we reveal that the solvability of the LME is equivalent to a mild structural condition, which is easily testable. The section concludes with a system of linear equations that can be used to recover a solution to the LME.
@@ IV-A MORP: Partition-Independent Sufficient Conditions
00 0 $=$ $C_{i}X_{i}+D_{i}U_{i}+F_{i}.$ (9)
Theorem 2 not only shows that the solvability of the partition-independent regulator equations (3) and LME (10) ensures the solvability of the regulator equations (3) but also provides a solution pair.
@@
$\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}Y_{i}=\begin{bmatrix}E_{i}\\ 0\end{bmatrix}$ (10)
@@
satisfies the regulator equations (3).
@@
where the second equation follows from $B_{i}Y_{i}=E_{i}$, the third equation is due to the fact that $A_{i}X_{i}+B_{i}U_{i}+E_{i}=X_{i}A_{0}$, and the fourth equation is a result of $\tilde{X}_{i}=s_{i}X_{i}$. By the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2), we further have
$C_{i}\tilde{X}_{i}+D_{i}\tilde{U}_{i}+s_{i}F_{i}$ $=s_{i}\left(C_{i}X_{i}+D_{i}(U_{i}+Y_{i})+F_{i}\right)-D_{i}Y_{i}$
$=s_{i}(C_{i}X_{i}+D_{i}U_{i}+F_{i})=0$ (13)
where the second equation follows from $D_{i}Y_{i}=0$ and the third equation is a consequence of $C_{i}X_{i}+D_{i}U_{i}+F_{i}=0$. We now conclude from (IV-A) and (IV-A) that the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2) satisfies the regulator equations (3). Hence the proof. ∎
@@
The converse of the first statement in Theorem 2 is not true. To show this, consider a MAS consisting of one leader and one follower with system parameters $A_{0}=0$, $A_{1}=B_{1}=C_{1}=D_{1}=1$, $E_{1}=-2$, $F_{1}=-1$ and $k$-partition transformation term $s_{1}=2$. It can be easily verified that the pair $(1,1)$ satisfies the regulator equations (3). We assume for contradiction that there is a pair $(X_{1},U_{1})$ that satisfies the partition-independent regulator equations (3) and a solution $Y_{1}$ to the LME (10). Then, (3) yields $X_{1}+U_{1}=2$ and $X_{1}+U_{1}=1$. Hence, $1=2$, a contradiction. We have just verified that the converse of the first statement in Theorem 2 is not true. One can also show that the LME (10) does not have a solution for the considered example.
@@
holds. Similar to the conditions in Theorem 2, this sufficient condition is not necessary for Condition 3 to hold. For, consider the example in Remark 8. What significantly distinguishes Theorem 2 from this sufficient condition is that Theorem 2 provides a solution pair, given by (2), to the regulator equations (3). This pair is explicitly expressed in terms of the $k$-partition transformation term and the matrices satisfying the partition-independent regulator equations (3) and LME (10). Therefore, in the second design strategy to be given, each follower can recompute its feedforward gain $K_{2i}$ without recomputing a solution to any regulator equations whenever the $k$-partition transformation term is updated.
@@ IV-B Discussion: Synthesis of Control Parameters
This subsection presents the second design strategy for the control parameters. Then, it compares both strategies from the perspectives of scalability and conservatism.
@@
The second design strategy involves solving the partition-independent regulator equations (3) and the LME (10). Note that neither of these equations includes the $k$-partition transformation term $s_{i}$. Therefore, once each follower finds a solution pair to the partition-independent regulator equations (3) and a solution to the LME (10), it can use these solutions to recompute the feedforward gain $K_{2i}$ whenever $s_{i}$ is changed. As a result, the design strategy in Remark 11 is much more scalable than the one in Remark 6.
@@
Theorem 2 establishes that if the second design strategy is applicable, so is the first design strategy. However, the converse is not true. To see this, consider the MAS described in Remark 8 and let $\mathcal{F}=1$. Observe that the conditions in Remark 6 are satisfied. However, we conclude from Remark 8 that neither step (iii) nor step (iv) in Remark 11 is feasible. Consequently, the first design strategy applies to a broader class of leader and follower dynamics.
@@ IV-C Solvability of the Introduced LME
This subsection starts with the straightforward characterizations of the solvability of the LME (10).
@@
    $\mathrm{im}\begin{bmatrix}E_{i}\\ 0\end{bmatrix}\subseteq\mathrm{im}\begin{bmatrix}B_{i}\\ D_{i}\\ \end{bmatrix}.$ (16)
@@
    $\mathrm{rank}\begin{bmatrix}B_{i}&E_{i}\\ D_{i}&0\end{bmatrix}=\mathrm{rank}\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}.$ (17)
We have the following immediate corollary, where (10) reduces to (18) for the follower dynamics not having a direct feedthrough term.
Corollary 2
Let $i\in\mathcal{N}$. Then the following conditions are equivalent:
  1. (i)
    There exists a solution $Y_{i}$ to the following LME
    $B_{i}Y_{i}=E_{i}.$ (18)
  2. (ii)
    The following inclusion holds:
@@
  3. (iii)
    The following rank condition holds:
    $\mathrm{rank}\begin{bmatrix}B_{i}&E_{i}\end{bmatrix}=\mathrm{rank}\hskip 0.85355ptB_{i}.$ (20)
Remark 14
The inclusion (16) and the inclusion (19) are the structural characterizations of the solvability of the LME (10) and the LME (18), respectively. They are easily testable by the rank conditions (17) and (20). The inclusion (19) is equivalent to follower $i$ being subject to only matched disturbances.
We now reformulate the LMEs (10) and (18) as systems of linear equations that can be used to recover solutions to them.
Remark 15
Let $i\in\mathcal{N}$. Suppose Condition 3 holds. By the proof of Theorem 1.9 in [32], a solution pair for the partition-independent regulator equations (3) can be obtained after transforming them into a system of linear equations of the form $Q_{i}z_{i}=b_{i}$ using the property (ii) of Proposition A.1 in [32]. Let the rank condition (17) hold. Similar to the partition-independent regulator equations case, a solution $Y_{i}$ to the LME (10) can be found by solving $Q_{i}z_{i}=b_{i}$ for $z_{i}$, where
$Q_{i}=I_{n_{0}}\otimes\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix},\ b_{i}=\mathrm{vec}\begin{bmatrix}E_{i}\\ 0\end{bmatrix},\ z_{i}=\mathrm{vec}\hskip 0.85355ptY_{i}$ (21)
and recovering $Y_{i}$ from $z_{i}$. A solution $Y_{i}$ to the LME (18) can be found by $Q_{i}z_{i}=b_{i}$ with $Q_{i}=I_{n_{0}}\otimes B_{i},\ b_{i}=\mathrm{vec}\hskip 0.85355ptE_{i},\ z_{i}=\mathrm{vec}\hskip 0.85355ptY_{i}$. Solutions to the LMEs (10) and (18) can also be obtained from Theorem 2 in [33] and computationally efficient iterative methods (e.g., see [34] and references therein).
V Illustrative Numerical Examples
Through three numerical examples, this section demonstrates the generality of the MORP and compares both design strategies regarding scalability. The following matrices are used to construct heterogeneous MASs in the examples
$A_{\alpha}$ $=$ $\begin{bmatrix}0.2&3\\ 0.1&-0.1\\ \end{bmatrix},\ B_{\alpha i}=\begin{bmatrix}0&4-i\\ 1&0\end{bmatrix},\ E_{\alpha i}=\begin{bmatrix}0&0\\ 0&i-5\end{bmatrix}$
$A_{\beta}$ $=$ $\begin{bmatrix}2&3&-1\\ 1&-1&0\\ 1&0&2\end{bmatrix},\ B_{\beta i}=\begin{bmatrix}i-5&0\\ i-3&0\\ 0&1\\ \end{bmatrix},\ E_{\beta i}=\begin{bmatrix}0&0\\ 0&0\\ i&0.5\end{bmatrix}$
$C$ $=$ $I_{2},\ D=\begin{bmatrix}0&0\\ 0&0.01\end{bmatrix},\ F=-I_{2},\ \Gamma=\begin{bmatrix}0&0.75\\ -0.75&0\end{bmatrix}.$
@@ V Illustrative Numerical Examples
[…]
Note that Conditions 2, 3, and 4 hold. Following the design steps in Remark 11, we first select $\mu=3$ to satisfy (8) and then design $K_{1i}$ using $lqr$ function in MATLAB [^3] For $i=1,\ldots 50$ ($i=51,\ldots,100$), take the state weight matrix as $15I_{2}$ ($25I_{3}$). For $i=1,\ldots,100$, take the input weight matrix as $I_{2}$.. As per steps (iii) and (iv) and Remark 15, for $i=1,\ldots,100$, a solution pair $(X_{i},U_{i})$ to the partition-independent regulator equations (3) and a solution $Y_{i}$ to the LME (18) are recovered from their equivalent systems of linear equations. Lastly, based on $k$-partition transformation of interest, we calculate $K_{2i}$ for each follower as suggested in step (v).
Since the bipartite framework partitions the followers [^4] There are studies incorporating the leader into the partition through the structural balance condition on $\mathcal{G}(\bar{\mathcal{A}})$ (e.g., see [17, 20]). This, however, allows only $1$ BORP to be solved. according to the signs of the entries of $\mathcal{A}$, for the considered $\mathcal{G}({\mathcal{A}})$, it yields a unique $2$-partition of the followers. Therefore, the existing formulations in [19, 18, 3] allow only $2$ BORPs to be solved by swapping the followers tracking the leader and its additive inverse. On the other hand, as discussed in Section II-A, the number of $2$-partitions of the followers and $1$-partition of the followers obtained by $k$-partition transformations are respectively $12^{99}-$ 1 and $1$, which are the corresponding Stirling numbers of the second kind. In fact, there are $2^{100}$ gauge transformations generating all the aforementioned $2^{99}$ partitions. Thus, the proposed formulation allows $2^{100}$ BORPs to be solved without changing the underlying graph.
In the simulation, we take the initial states as $x_{i}^{T}(0)=[3i/50,0]$ for $i=1,\ldots,50$, $x_{i}^{T}(0)=[-3i/50+3,0,0]$ for $i=51,\ldots,100$, $\eta_{i}(0)=0$ for $i=1,\ldots,100$, and $v(0)=1$. The top row of Fig. 1 illustrates the output responses of 2 BORPs that can be solved using the bipartite framework and the proposed formulation. For these BORPs, the design [^5] There is a typo in the equation (4b) of [19]. For Theorem 1 in [19] to be valid, the term $(z_{j}-\mathrm{sgn}(a_{ij})z_{i})$ needs to be replaced with $(\mathrm{sgn}(a_{ij})z_{j}-z_{i})$. in [19] and our design generate identical output responses. The bottom row of Fig. 1 presents $2$ out of $22^{100}-$ 2 BORPs that can be solved with the proposed formulation but not with the existing bipartite framework.
Figure 1: The top row depicts the identical output responses with the design in [19] and the proposed design. The bottom row presents the output responses with the proposed formulation for 2 BORPs that are impossible to formulate with the approach in [19] without changing the underlying graph.

### v2

@@ III-A Distributed Control Law
Any distributed control law solving the CORP for the MAS composed of (III-A) and (2) over an augmented unsigned directed graph solves the MORP over the same graph.
In the light of Lemma 2, we consider the following feedforward-based distributed control law [^1] With the distributed measurement output feedback control law in Equation (8.14) of [1], one can arrive at a result similar to Theorem 1 (ii) under the additional detectability assumptions. The resulting design strategy will have the drawback in Remark 7. In this case, Theorem 2 is still a remedy. proposed in [5]
@@
where $\eta_{i}\in\mathbb{R}^{n_{0}}$ is the estimate of $v$. The state equation in (III-A) is called a distributed observer. Moreover, $\mu\in\mathbb{R}$, $K_{1i}\in\mathbb{R}^{m_{i}\times n_{i}}$, and $K_{2i}\in\mathbb{R}^{m_{i}\times n_{0}}$ are control parameters to be designed. Compared to the distributed observer in [5], the one in (III-A) uses the absolute value of the adjacency matrix entries to be readily applicable even if the given graph is signed.
@@ III-B MORP: Necessary and Sufficient Conditions
The augmented signed directed graph $\mathcal{G}(\bar{\mathcal{A}})$ contains a directed spanning tree [^2] By the definition of the augmented signed directed graph given in Section II-B, the root of any directed spanning tree in $\mathcal{G}(\bar{\mathcal{A}})$ is necessarily node 00..
@@
The $k$-partition transformation term $s_{i}$ is incorporated into (3) due to Lemma 2. Except for this term, Condition 3 is standard for the studies investigating the CORP with the feedforward approach (e.g., see [5, 8]). The articles [17, 15] studying the BORP with the feedforward approach consider the disturbance-free (i.e., $E_{i}=0$ and $G_{i}=0$) and direct feedthrough-free (i.e., $D_{i}=0$) follower dynamics. However, introducing a gauge transformation term in the regulator equations for each follower can extend their distributed controllers to take the effect of disturbances and direct feedthrough into account. To avoid such modification, the authors in [16] assume that agents in different sets in partition are subject to disturbances that are additive inverses of each other. This can be impractical in real-world applications, for instance, consider networked UAVs operating in the same environment. Thus, this article does not make that assumption about disturbances.
@@ III-C Discussion: Synthesis of Control Parameters
This subsection presents the first design strategy for the control parameters and highlights an associated drawback. Let $\mathcal{L}_{u}$ denote the Laplacian matrix of the unsigned directed graph $\mathcal{G(|\mathcal{A}|)}$ and $\mathcal{H}=\mathcal{L}_{u}+\mathcal{F}$. Let $\lambda_{i}(\mathcal{H})$ and $\lambda_{j}(A_{0})$ denote the eigenvalues of $\mathcal{H}$ and $A_{0}$ for $i=1,\ldots,N$, $j=1,\ldots,n_{0}$.
@@
    Select $\mu$ according to the inequality333 The sufficient and necessary conditions for the Hurwitzness of the matrix $A_{\mu}=(I_{N}\otimes A_{0})-\mu(\mathcal{H}\otimes I_{n_{0}})$, which is the system matrix of the distributed observer in compact form, are used in the proof of Theorem 1. These conditions are given in Lemma 4 of [30]. However, the lower bound on $\mu$ in (8), due to Lemma 3 given in Appendix, has no conservatism compared to the bound in Lemma 4 of [30]. Another lower bound is also provided in Lemma 3.2 of [1]. But, one can easily construct a counterexample to the first statement of that lemma by considering a Hurwitz $A_{0}$.
@@
As $\mathrm{spec}(\mathcal{H})\subsetneq\mathrm{ORHP}$ under Condition 4 (e.g., see Lemma 1 in [5]), any positive $\mu$ satisfies the inequality (8) if $\mathrm{spec}(A_{0})\subsetneq\mathrm{CLHP}$. Thus, the design of $\mu$ is independent of the eigenvalues of $\mathcal{H}$ (i.e., the spectral property of $\mathcal{G}(\bar{\mathcal{A}})$) when the exosystem generates a linear combination of constant signals, sinusoidal signals, polynomial signals, for instance, ramp signals, and exponentially converging signals. These signals cover references and disturbances encountered in most multi-agent systems (e.g., see the applications in [3, 6, 9, 31]).
@@
Though the design based on Remark 6 is straightforward to apply, it has a drawback in that the pair $(X_{i},U_{i})$ for each follower depends on $s_{i}$. Therefore, whenever the $k$-partition transformation is updated, it necessitates each follower to recompute a solution pair for the regulator equations (3) unless its $s_{i}$ remains unchanged. Hence, given a finite set of $k$-partition transformation terms with $M$ elements, each follower computes $M$ solution pairs to (3). This drawback can render the design strategy impractical in some applications. For example, consider low-cost networked UAVs trying to avoid radars. In this scenario, the set of $k$-partition transformation terms for each follower may not be known before the operation. Thus, it may be desirable in terms of computational cost to find a solution pair for the regulator equations (3) once and use it throughout the operation.
@@ IV MORP: Partition-Independent Solvability
This section first provides the partition-independent sufficient conditions for the solvability of the MORP. For each $i\in\mathcal{N}$, these conditions include the partition-independent regulator equations, as in the CORP with the feedforward approach, and an LME depending only on $B_{i}$, $E_{i}$, $D_{i}$, and $G_{i}$. The constructive nature of the result yields the second design strategy that eliminates the drawback discussed in Remark 7. Lastly, we reveal that the intoduced LME’s solvability is equivalent to an easily testable mild structural condition.
@@ IV-A MORP: Partition-Independent Sufficient Conditions
00 0 $=$ $C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i}.$ (9)
Theorem 2 not only shows that the solvability of the partition-independent regulator equations (3) and the LME (10) ensures the solvability of the partition-dependent regulator equations (3) but also provides a solution pair.
@@
$\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}Y_{i}=\begin{bmatrix}E_{i}\\ G_{i}\end{bmatrix}$ (10)
@@
satisfies the partition-dependent regulator equations (3).
@@
where the second equation follows from $B_{i}Y_{i}=E_{i}$, the third equation is due to the fact that $A_{i}X_{i}+B_{i}U_{i}+E_{i}=X_{i}A_{0}$, and the fourth equation is a consequence of $\tilde{X}_{i}=s_{i}X_{i}$. By the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2), we further have
$\hskip 2.84544ptC_{i}\tilde{X}_{i}+D_{i}\tilde{U}_{i}+G_{i}+s_{i}F_{i}$
$=$ $\hskip 2.84544pts_{i}\left(C_{i}X_{i}+D_{i}(U_{i}+Y_{i})+F_{i}\right)-D_{i}Y_{i}+G_{i}$
$=$ $\hskip 2.84544pts_{i}(C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i})=0$ (13)
where the second equation follows from $D_{i}Y_{i}=G_{i}$ and the third equation is a result of $C_{i}X_{i}+D_{i}U_{i}+G_{i}+F_{i}=0$. We now conclude from (IV-A) and (IV-A) that the pair $(\tilde{X}_{i},\tilde{U}_{i})$ in (2) satisfies the partition-dependent regulator equations (3). Hence, the proof is over. ∎
@@
The converse of the first statement in Theorem 2 is not true. To show this, consider a MAS consisting of one leader and one follower with system parameters $A_{0}=G_{1}=0$, $A_{1}=B_{1}=C_{1}=D_{1}=1$, $E_{1}=-2$, $F_{1}=-1$, and $k$-partition transformation term $s_{1}=2$. It can be verified that the pair $(1,1)$ satisfies the partition-dependent regulator equations (3). We assume for contradiction that there is a pair $(X_{1},U_{1})$ that satisfies the partition-independent regulator equations (3) and a solution $Y_{1}$ to the LME (10). Then, (3) yields $X_{1}+U_{1}=2$ and $X_{1}+U_{1}=1$. Hence, $1=2$, a contradiction. We have just proved that the converse of the first statement in Theorem 2 is not true. One can also show that the LME (10) does not have a solution for the considered example.
@@
holds. Similar to the conditions in Theorem 2, this sufficient condition is not necessary for Condition 3 to hold. To see this, consider the example in Remark 8. What significantly distinguishes Theorem 2 from this sufficient condition is that Theorem 2 provides a solution pair, given by (2), to the partition-dependent regulator equations (3). This pair is explicitly expressed in terms of the $k$-partition transformation term and the matrices satisfying the partition-independent regulator equations (3) and the LME (10). Thus, in the second design strategy to be given, each follower can recompute its feedforward gain $K_{2i}$ without recomputing a solution to any equations when the $k$-partition transformation term is updated.
@@ IV-B Discussion: Synthesis of Control Parameters
This subsection first presents the second design strategy for the control parameters. Then, it compares both strategies from the perspectives of scalability and conservatism.
@@
The second design strategy involves solving the partition-independent regulator equations (3) and the LME (10). Neither of these equations includes the $k$-partition transformation term $s_{i}$. Therefore, once each follower finds solutions to the partition-independent regulator equations (3) and the LME (10), it can use these solutions to recompute the feedforward gain $K_{2i}$ whenever $s_{i}$ is changed. As a result, the design strategy in Remark 11 is much more scalable than the one in Remark 6, as illustrated in Example 2.
@@
Theorem 2 establishes that if the second design strategy is applicable, so is the first design strategy. Yet, the converse is not true. To see this, consider the MAS described in Remark 8 and let $\mathcal{F}=1$. Observe that the conditions in Remark 6 are satisfied. However, we conclude from Remark 8 that neither step (iii) nor step (iv) in Remark 11 is feasible. Consequently, the first design strategy applies to a broader class of leader and follower dynamics.
Lastly, Table I summarizes the differences between the first and second design strategies.
TABLE I: Differences in Design Strategies
Differences Strategies First Second     Conclusion
[…]
LMEs to be solved Partition-Dependent: (3) Partition-Independent:         (3) and (10)     The second is more scalable, as discussed in Remark 12.
@@ IV-C Solvability of the Introduced LME
This subsection provides the straightforward characterizations of the solvability of the LME (10).
@@
    $\mathrm{im}\begin{bmatrix}E_{i}\\ G_{i}\end{bmatrix}\subseteq\mathrm{im}\begin{bmatrix}B_{i}\\ D_{i}\\ \end{bmatrix}.$ (16)
@@
    $\mathrm{rank}\begin{bmatrix}B_{i}&E_{i}\\ D_{i}&G_{i}\end{bmatrix}=\mathrm{rank}\begin{bmatrix}B_{i}\\ D_{i}\end{bmatrix}.$ (17)
Remark 14
The inclusion (16) is a structural characterization of the solvability of the LME (10). It is also easily testable by the rank condition (17). The inclusion (16) holds only when follower $i$ is subject to solely matched disturbances, which are common in certain applications (e.g., nonholonomic wheeled robots [33], UAVs [34], and spacecrafts [35]). Nevertheless, the first design strategy can still be employed in the presence of unmatched disturbances.
V Experimental and Numerical Illustrations
This section demonstrates the MORP’s flexibility in shifting mission objectives via an experiment with networked mobile robots. It also provides two numerical examples to showcase the MORP’s generality and compare the first and second design strategies regarding scalability. The following matrices are used throughout this section for the dynamics of the MASs.
$A_{\alpha}$ $=$ $\begin{bmatrix}0_{2\times 2}&I_{2}\\ 0_{2\times 2}&0_{2\times 2}\end{bmatrix},\ \Gamma=\begin{bmatrix}0&0.0025\\ -0.0025&0\end{bmatrix},$
$A_{\beta}$ $=$ $\begin{bmatrix}0.2&3\\ 0.1&-0.1\\ \end{bmatrix},\ B_{\beta}=\begin{bmatrix}0&3\\ 1&0\end{bmatrix}.$
Experiment 1
@@ V Experimental and Numerical Illustrations
We assume that cylindrical and cuboid obstacles in the operating terrain are detected through a distributed sensor fusion algorithm that runs onboard each robot. Accordingly, the robots update their $k$-partition transformation terms. To simulate this, the $k$-partition transformation is defined as a piecewise constant function. Specifically, $S(t)=\mathrm{diag}(1,0.75,0.5)$ for $t\in[0,86.5)\cup[161,212]$ seconds and $S(t)=\mathrm{diag}(2.3,1.65,1)$ for $t\in[86.5,161)$ seconds. Note that Conditions 2, 3, 4, and the inclusion (16) hold. Following Remark 11, we set $\mu=10$ and design $K_{1i}$ using $place$ function in MATLAB [^5] For $i=1,2,3$, $\mathrm{spec}(A_{i}+B_{i}K_{1i})=\{-0.75,-1.25,-1.75,-2.5\}$.. As per steps (iii) and (iv), for $i=1,2,3$, a solution pair $(X_{i},U_{i})$ to the partition-independent regulator equations (3) is recovered from their equivalent system of linear equations (see the proof of Theorem 1.9 in [32]) and a solution $Y_{i}$ to the LME (10) is found using $linsolve$ function in MATLAB [^6] The authors thank Jackson Kulik for his comment on solving (10).. Lastly, based on $S$, we calculate $K_{2i}$ for each follower as in step (v).
The experiment is initiated with $x_{1}^{\mathrm{T}}(0)=[1.2,1.5,0,0]$, $x_{2}^{\mathrm{T}}(0)=[0.1,1.7,0,0]$, $x_{3}^{\mathrm{T}}(0)=[-0.5,1.3,0,0]$, $\eta_{i}(0)=0$ for $i=1,2,3$, and $v^{\mathrm{T}}(0)=[0,1]$. As seen in Fig. 1, the output of follower $i$ tracks the $s_{i}$ multiple of the leader’s state and steers around obstacles successfully for $i=1,2,3$. The experiment video can be found in the clickable link.
Figure 1: The trajectories of the MAS. Here, $y_{ij}$ denotes the $j$th entry of follower $i$’s output for $i=1,2,3$ and $y_{0j}$ denotes the $j$th entry of $v$ while “$\times$” and “o” marks $y_{ij}$ at the initial and final times for $i=0,1,2,3$, respectively.
@@
As indicated in Remark 2, the MORP includes the BORP. To make the MORP comparable with the existing solutions to the BORP, we force $k$-partition transformations to be gauge transformations. Despite such a restriction, this example presents the generality of the MORP. To this end, consider $100$ followers and a leader with the following matrices: $A_{i}=A_{\beta}$, $B_{i}=B_{\beta}$, $C_{i}=[1,0]$ for $i=1,\ldots,50$; $A_{i}=A_{\alpha}$, $B^{T}_{i}=[0\ I_{2}]$, $C_{i}=[1,0,0,0]$ for $i=51,\ldots,100$; $D_{i}=0$, $E_{i}=0$, $G_{i}=0$, $F_{i}=-1$ for $i=1,\ldots,100$; $A_{0}=0$. They communicate over $\mathcal{G}(\bar{\mathcal{A}})$ with $a_{i1}=1$ for $i=3,5,\ldots,99$, $a_{i1}=-1$ for $i=2,4,\ldots,100$, and $f_{1}=1$, while the remaining entries of $\mathcal{A}$ and $\mathcal{F}$ are zero.
As the bipartite framework partitions the followers [^7] There are studies incorporating the leader into the partition through the structural balance condition on $\mathcal{G}(\bar{\mathcal{A}})$ (e.g., see [15]). This, however, allows only $1$ BORP to be solved. based on the signs of the entries of $\mathcal{A}$, for the considered $\mathcal{G}({\mathcal{A}})$, it yields a unique $2$-partition of the followers. Therefore, the existing formulations in [17, 16, 18] allow only $2$ BORPs to be solved by swapping the followers that track the leader’s state and its additive inverse. On the other hand, as discussed in Section II-A, the number of $2$-partitions of the followers and $1$-partition of the followers obtained by $k$-partition transformations are respectively $12^{99}-$ 1 and $1$, which are the corresponding Stirling numbers of the second kind. In fact, there are $2^{100}$ gauge transformations generating all the aforementioned $2^{99}$ partitions. Thus, the proposed formulation allows $2^{100}$ BORPs to be solved without changing the underlying graph.
The simulation is initiated with $x_{i}^{\mathrm{T}}(0)=[3i/50,0]$ for $i=1,\ldots,50$, $x_{i}^{\mathrm{T}}(0)=[-3i/50+3,0,0,0]$, for $i=51,\ldots,100$, $\eta_{i}(0)=0$ for $i=1,\ldots,100$, and $v(0)=1$. The top row of Fig. 2 illustrates the output responses of 2 BORPs that can be solved using the bipartite framework and the proposed formulation, where $S=\pm\left(I_{50}\otimes\mathrm{diag}(1,-1)\right)$. For these BORPs, the design [^8] There is a typo in Equation (4b) of [17]. For Theorem 1 in [17] to be valid, the term $(z_{j}-\mathrm{sgn}(a_{ij})z_{i})$ needs to be replaced with $(\mathrm{sgn}(a_{ij})z_{j}-z_{i})$. in [17] and the first and second design strategies can generate identical output responses. The bottom row of Fig. 2 presents $2$ out of $22^{100}-$ 2 BORPs that can be solved with the proposed formulation but not with the bipartite framework under the same graph. Here, $S=\pm\left(\mathrm{diag}(1,-1)\otimes I_{50}\right)$.
Figure 2: The top row depicts the identical output responses with the design in [17] and the proposed design. The bottom row presents the output responses with the proposed formulation for 2 BORPs that are impossible to formulate with the approach in [17] without altering the underlying graph.